One has the following information about the sector of companies: Direct tax: 10% of the Profits from production and patrimony (income), Distributed Profits: 117% of the Direct tax, Current transfers toward other sectors: 6% from the Disposable Income, Disposable income: 56.9 billion Euro, Profits from production and patrimony (income): 74.4 billion Euro. Use an account of the sector of companies and compute the value of the current transfers received from other sectors!


a.The value of the current transfers received from other sectors is 25.1 billion Euro

b.The value of the current transfers received from other sectors is 14.6 billion Euro

c.The value of the current transfers received from other sectors is 5.6 billion Euro

d.The value of the current transfers received from other sectors is 1.0 billion Euro

e.The value of the current transfers received from other sectors is 2.1 billion Euro




1
Expert's answer
2022-01-23T15:40:15-0500

"CTR=p(b-a)\frac n{1-\eta}=14.6" billion Euro.